OpenStudy (anonymous):

Use the rules of exponents to evaluate or simplify. Write without negative exponents. (-243)^(3/5)

OpenStudy (solomonzelman):

\[\huge\color{purple}{a^{b/c}=\sqrt[c]{a^b}}\]

OpenStudy (solomonzelman):

Sorry got disconnected.

OpenStudy (solomonzelman):

\[\huge\color{purple}{(-243)^{3/5}}\] \[\huge\color{purple}{\sqrt[~~5]{(-243)^3}}\] \[\huge\color{purple}{\sqrt[~~5]{(-3^5)^3}}\] \[\huge\color{purple}{\sqrt[~~5]{(-3)^{(3 \times 5)}}}\] \[\huge\color{purple}{\sqrt[~~5]{((-3)^3)~^5}}\] \[\huge\color{purple}{\sqrt[~~5]{(-27)^5}}\] \[\huge\color{purple}{-27}\]